OpenStudy (anonymous):

Determine and simplify dy/dx:

OpenStudy (anonymous):

cancel the \(d\)'s top and bottom and get \[\frac{y}{x}\]

OpenStudy (anonymous):

can't read it sorry

OpenStudy (anonymous):

\[y=tanx/xcosx\]

OpenStudy (anonymous):

\[y=x ^{7}e ^{x ^{7}}+8^{x-1}\times4^{-x ^{2}-secx ^{3}}\]

OpenStudy (anonymous):

first one requires the quotient rule , also the product rule \[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\]with \[f(x)=\tan(x), f'(x)=\sec^2(x), g(x)=x\cos(x), g'(x)=\cos(x)-x\sin(x)\]

OpenStudy (anonymous):

second one is going to be a real pita the first part is not too bad, \[(fg)'=f'g+g'f\] with \[f(x)=x^7, f'(x)=7x^6, g(x)=e^{x^7}, g'(x)=7x^6e^{x^7}\]

OpenStudy (anonymous):

second part is going to be just a total mess \[\frac{d}{dx}8^{x-1}=\ln(8)8^{x-1}\]

OpenStudy (anonymous):

i think @FibonacciChick666 will finish finding the derivative of \[\large 4^{-x^2-\sec(x^3)}\]

OpenStudy (anonymous):

who makes up these functions?

OpenStudy (fibonaccichick666):

... you have to be kidding me... that's just evil

OpenStudy (fibonaccichick666):

alright, so first, let's clean up the given

OpenStudy (fibonaccichick666):

that isn't multiplication of the x^2 is it...

OpenStudy (fibonaccichick666):

oh yuck

OpenStudy (fibonaccichick666):

and actually @satellite73 it's this: \[y=x ^{7}e ^{x ^{7}}+8^{x-1}.4^{-x ^{2}-Secx ^{3}}\]according to the word document

OpenStudy (anonymous):

yes, please go with the word document question

OpenStudy (fibonaccichick666):

so just the \(.4^{-x^2-secx^3}\) for me. first we need to substitute in for the exponent so let's set \(g=-x^2-secx^3\) so all you have to do @Jas1997 is find the derivative of \(.4^g\) then what the derivative of g is for the second part of the product rule.

OpenStudy (bradely):

step by step solution http://www.mathskey.com/question2answer/17510/derivative

OpenStudy (anonymous):

Thankyou @bradely @FibonacciChick666 @satellite73